a repair company's charge for repairing a certain type of copy machine fits the model y=47.38 + 0.617x where y is the number of dollars charged and x is the number of minutes the repair person the job. How many minutes would it take for the cost of repair to reach $60
Answer by Cromlix(4381)

Hi there,
y=47.38 + 0.617x
47.38 + 0.617x = 60
0.617x = 60 - 47.38
0.617x = 12.62
x = 12.62/0.617
x = 20.45 minutes (2 decimal places)
Hope this helps :-)